Prepare a table showing ten years of payments on a 10-year loan of $1,000,000 at annual interest = 5%. Include all information shown on the following sample table.

Year Loan Payment Loan Interest Principal Payment Balance remaining
0 1,000,000
1
2

Revise the table with a new calculation that results from an extra payment of $200,000 at the end of year 3.
Answer #1

Principal Loan = $1,000,000

Interest = 5%

Tenure = 10 Years

In order to create the amortization schedule, below steps must be followed:

1. For year 1, Start with Opening principal amount, note it.

2. Compute the Yearly Loan payment using the MS Excel function PMT

= PMT (rate, nper, pv, fv, [type])

= PMT (0.05, 10, -1000000, 0)

= $129,504.57

3. Compute the Loan Interest on Opening Principal amount using simple interest formula,

Loan Interest = (Principal x rate x Time) / 100

This is the interest paid for the year.

4. Subtract the Interest amount from yearly instalment amount in order to compute Principal Payment.

5. Subtract the Principal paid from Opening principal to compute Balance Remaining.

6. For next year, the closing principal of previous year is the opening principal of current year.

7. Repeat steps 2 to 6 till Year 10.

Below table gives the various values and computation for the repayment of loan.

Yearly Instalment 129504.57 =PMT(0.05,10,-1000000,0)
Year Opening principal Loan Interest Loan Payment Principal Payment Closing Principal
1 1000000 50000 129505 79505 920495
2 920495 46025 129505 83480 837016
3 837016 41851 129505 87654 749362
4 749362 37468 129505 92036 657325
5 657325 32866 129505 96638 560687
6 560687 28034 129505 101470 459217
7 459217 22961 129505 106544 352673
8 352673 17634 129505 111871 240802
9 240802 12040 129505 117464 123338
10 123338 6167 129505 123338 0

For the case when extra payment of $200,000 is made at the end of year 3 (beginning of year 4), the revised table with a new calculation is as below:

Yearly Instalment 129504.57 =PMT(0.05,10,-1000000,0)
Year Opening principal Loan Interest Loan Payment Principal Payment Closing Principal
1 1000000 50000 129505 79505 920495
2 920495 46025 129505 83480 837016
3 837016 41851 129505 87654 749362
4 749362 37468 329505 292036 457325
5 457325 22866 129505 106638 350687
6 350687 17534 129505 111970 238717
7 238717 11936 129505 117569 121148
8 121148 6057 129505 123447 -2299
9 -2299 -115 129505 129620 -131919
10 -131919 -6596 129505 136101 -268019

The loan is completely repaid in Year 8.